OpenStudy (anonymous):

2. (2x2 – 23x + 63) ÷ (x – 7) use synthetic division. Show work for full credit.

OpenStudy (ash2326):

(2x^2-23x+63)/ x-7 7| 2 -23 63 14 -63 2 -9 0 so (2x^2-23x+63)/(x-7)= 2x-9

OpenStudy (anonymous):

thank u do u think u cud help me with 2 more?

OpenStudy (ash2326):

yeah post them

OpenStudy (anonymous):

3. Let f(x) = 4x3 + 7x2 – 13x – 3 and g(x) = x + 3. Find

OpenStudy (anonymous):

find f(x)/g(x)

OpenStudy (ash2326):

(4x^3+7x^2-13x-3)/(x+3) -3 | 4 7 -13 -3 -12 15 -6 4 -5 2 -9 so we'll get (4x^2-5x+2)+ -9/(x+3)

OpenStudy (anonymous):

4. Let f(x) = 3x3 – 4x – 1 and g(x) = x + 1. find f(x)/g(x)

OpenStudy (ash2326):

(3x^3-4x-1)/(x+1) -1 | 3 0 -4 -1 -3 +3 1 3 -3 -1 0 so f(x)/g(x)= 3x^2-3x-1

OpenStudy (anonymous):

thank u so much ash

OpenStudy (ash2326):

welcome:)

OpenStudy (anonymous):

ash you think you can help help me with the same type of exercice?

OpenStudy (ash2326):

yeah post it as a new question. I'll help you:)

OpenStudy (anonymous):

Let f(x) = x4 – 8x3 + 16x2 – 19 and g(x) = x – 5. Find f(x)/g(x)

OpenStudy (ash2326):

@ricky10messi10 please post it as a new question:)